# Hyper-STTN: Social Group-aware Spatial-Temporal Transformer Network for Human Trajectory Prediction with Hypergraph Reasoning

### v1

@@ Abstract
Predicting crowded intents and trajectories is crucial in varouls real-world applications, including service robots and autonomous vehicles. Understanding environmental dynamics is challenging, not only due to the complexities of modeling pair-wise spatial and temporal interactions but also the diverse influence of group-wise interactions. To decode the comprehensive pair-wise and group-wise interactions in crowded scenarios, we introduce Hyper-STTN, a Hypergraph-based Spatial-Temporal Transformer Network for crowd trajectory prediction. In Hyper-STTN, crowded group-wise correlations are constructed using a set of multi-scale hypergraphs with varying group sizes, captured through random-walk robability-based hypergraph spectral convolution. Additionally, a spatial-temporal transformer is adapted to capture pedestrians’ pair-wise latent interactions in spatial-temporal dimensions. These heterogeneous group-wise and pair-wise are then fused and aligned though a multimodal transformer network. Hyper-STTN outperformes other state-of-the-art baselines and ablation models on 5 real-world pedestrian motion datasets.
@@ I Introduction
Human trajectory prediction is a pivotal area in computer vision and robotics research, focusing on predicting agents’ future movements by analyzing their past behaviors. This is vital for many real-world applications such as social robot navigation [1, 2] and smart city systems [3]. Despite its importance, forecasting human trajectories in social settings is complex due to the unpredictable nature and varied patterns of crowd movements. The challenge is intensified by the need to understand the intricate interactions and relations among different social groups, which involve complex cooperation and competitive dynamics, as shown in Fig. 1.
@@
Figure 1: HHI (human-human interaction) illustration: Group-wise HHI captures latent correlations among high-level perspectives on group behaviors, while pair-wise spatial-temporal dependencies represent individual influences.
@@
To address these challenges, we introduce Hyper-STTN, a hypergraph-based spatial-temporal transformer network designed to handle both pair-wise and group-wise underlying social interactions, as illustrated in Fig. 1, across spatial-temporal dimensions. The core concept behind Hyper-STTN is its utilization of an effective HHI and social interaction inference framework using multiscale hypergraphs. The main contributions of this paper can be summarized as follows:
  • We propose a novel human trajectory prediction framework called Hyper-STTN, which constructs crowded dynamics using a set of multiscale hypergraphs to leverage group-wise and pair-wise social interactions in both spatial and temporal dimensions.
  • Hyper-STTN aligns heterogeneous features and interactions using a multi-modal transformer network to avoid confusion in interpretations.
  • Hyper-STTN demonstrates outstanding performance compared with other existing SOTA approaches on publicly available pedestrian trajectory datasets.
@@ II-A Related Works
Earlier efforts formulated the task of predicting human trajectories based on handcrafted models, such as the social force model [11] and Gaussian process [12], which utilized fixed physical or mathematical rules to adapt to environmental dynamics. The inflexibility of traditional approaches could lead to overlooking underlying social interaction and forecasting errors in more complex conditions. Inspired by the advancements in machine learning, many deep learning-based algorithms have been deployed to better understand social interactions among pedestrians [5, 7]. However, previous learning-based methods either could not fully utilize the captured features directly or could not effectively leverage long-term dependencies. Especially, due to the successes of attention mechanisms and transformers in sequence learning and pair-wise feature representation [13], [8, 14], there have been explicit descriptions of social interactions via both spatial and temporal dimensions using a spatiotemporal graph model of crowd movements. While the description of pair-wise interactions has been well addressed with the development of transformers and attention modules, all the aforementioned methods have neglected to collect group-wise latent influences for individuals.
While [6] has adapted a hypergraph-based network to learn group-wise and pair-wise features with the multiscale property of hypergraph generation, GroupNet still struggles with isolating spatial features along the temporal dimension. Therefore, the challenge of aligning both group-wise and pair-wise social interactions with long-term dependencies remains a key problem in human trajectory prediction. In this work, we not only incorporate spatial and temporal pair-wise interactions into hypergroup-based group-wise interaction representation, but also introduce a cross-modal attention mechanism to leverage the heterogeneous multimodality dependencies mentioned aboive, based on a multi-modal transformer [15].
@@ III-A Preliminary
We adapt a set of multiscale hypergraphs $\mathcal{G}=(\mathcal{V,E,T}\mathbf{,W})$ to construct crowd movement dynamics for trajectory prediction inference. Where $\mathcal{V}=\{{v_{\rm 1}},{v_{\rm 2}},\cdots,{v_{\rm N}}\}$ is the set of vertexes that presents total ${\rm N}$ agents in the scenario, and =$\mathcal{E}=\{{e_{\rm 1}},{e_{\rm 2}},\cdots,{e_{\rm M}}\}$ denotes that a set of ${\rm M}$ hyperedges are constructed in the hypergraphs, particularly each hyperedge can link more than two vertexes to form a group, and $\mathcal{T}=\{{\tau_{\rm 1}},{\tau_{\rm 2}},\cdots,{\tau_{\rm H}}\}$ presents $\rm H$ scales of hypergrpahs, and $\mathbf{W}=diag({{w}_{\rm e_{1}}},{{w}_{\rm e_{2}}},\cdots,{{w}_{\rm e_{M}}})\in\mathbb{R}^{M\times M}$ defines a weighted matrix of hypergraphs. Lastly, the incident matrix $\mathbf{H}\in\mathbb{R}^{N\times M}$ of hypergraph $\mathcal{G}$ is defined as follows: $\mathbf{H}(v,e)=1,~{}\text{if}~{}v\in e;~{}\mathbf{H}(v,e)=0,~{}\text{if}~{}v\notin e.$ The diagonal vertex matrix $\mathbf{M}_{v}\in\mathbb{R}^{N\times N}$ and diagonal edge matrix $\mathbf{M}_{e},\mathbf{W}\in\mathbb{R}^{M\times M}$ are composed by the degree of vertex $d(v)=\sum_{e\in\mathcal{E}}w_{(e)}\mathbf{H}(v,e)$ and the degree of edge $d(e)=\sum_{v\in\mathcal{V}}\mathbf{H}(v,e)$.
Additionally, the adjacency matrix of hypergraph $\mathcal{G}$ can be represented by $\mathbf{A}\in\mathbb{R}^{N\times N}$, whereas $\mathbf{A}=\mathbf{H}\mathbf{W}\mathbf{H}^{\top}-\mathbf{M}_{\rm v}$. Hyper-STTN not only models group-wise interaction using pedestrians hypergraph construction, but also leverages agents’ pair-wise spatial and temporal social interaction into the hypergraph from spatial-temporal transformer. Let $\mathbf{X}\in\mathbb{R}^{N\times T_{i}\times d}$ and $\mathbf{\hat{X}}\in\mathbb{R}^{N\times T_{o}\times d}$ denote the past and future trajectory data of crowd with time sequence $T_{i}=8$ or $T_{o}=12$ and dimension $d=2$. And the position of $i$th agent in time $t$ can be defined as $\mathbf{x}_{i}^{t}=(x,y)$. Lastly, Hyper-STTN describes the pedestrian trajectory distribution $\mathcal{P}$ as follows:
@@ III-B Hyper-STTN Architecture
Hyper-STTN captures group-wise and pair-wise interactions to reason out HHI for human trajectory prediction tasks, as illustrated in Fig. 2. Firstly, the spatial-temporal transformer network is developed to capture spatial and temporal dimensional pair-wise features. Meanwhile, the hypergraph convolution neural network is employed in parallel to present crowd group-level dependencies, based on crowded multiscale hypergraph. Subsequently, above heterogeneous group-wise and pair-wise crowded interaction matrices are fused through a multi-modal transformer to align the spatial-temporal embeddings into the hypergraph. Finally, the crowd embedding mixture is decoded from a CVAE decoder to forecast future trajectories.
@@ III-C Spatial-Temporal Transformer
Inspired by [3, 8, 1], we have developed a spatial-temporal transformer network to capture the correlation of agents’ movement along spatial-temporal dimensions. Particularly, spatial or temporal dependencies can be represented as multiple attention maps, as shown in Fig. 6. In the spatial transformer, the input data $\mathbf{X}$ is first fed into a positional encoding (PE) layer for sequence information encoding based on [13]. Subsequently, the input data proceeds in the proper order through a layer normalization block (LN), a multihead spatial attention layer, and a feed-forward network (FFN), as shown in Fig. 4. Additionally, residual connections are deployed to stabilize the network. The core of the spatial-temporal transformer is the attention mechanism, where the calculation of the multihead spatial or temporal attention layer is defined as follows:
@@
where $\operatorname{Atten}(\cdot)$ is self-attention mechanism function, $\operatorname{Multi}$ is multihead self-attention mechanism function, and $\rm{head}_{(\cdot)}$ denotes $\rm i$-th head in the multihead attention function with $\rm{i}\in[1,\cdots,\rm{h}]$. $f_{\rm{fc}}$ is a fully connected layer, $\rm\mathbf{Q}^{i},\rm\mathbf{K}^{i},\rm\mathbf{V}^{i}$ are $\rm i$-th query matrix, key matrix, and value matrix with the dimension $\rm d_{h}$.
Figure 3: Group-wise HHI Representation: i) We construct group-wise HHI with a set of multiscale hypergraphs, where each agent is queried in the feature space with varying ‘k’ in KNN to link multiscale hyperedges. ii) After constructing HHI hypergraphs, group-wise dependencies are captured by point-to-edge and edge-to-point phases with hypergraph spectral convolution operations.
@@
The framework of temporal transformer and spatial transformer is illustrated as Fig. 4. And spatial transformer captures the pair-wise spatial interaction $\mathbf{\hat{Y}}_{\rm{S}}\in\mathbb{R}^{N\times T_{i}\times d}$ of all agents along each timestep. In contrary, temporal transformer rotates the input dimensions to express individual’s temporal interaction $\mathbf{\hat{Y}}_{\rm{T}}\in\mathbb{R}^{N\times T_{i}\times d}$ along each agent.
@@ III-D Hypergraph Neural Network
To couple multiple group sizes of crowd, we generate a set of multiscale hypergraphs to present social group classifications for comprehensive intergraph or intragraph interactions representation via hypergraph convolution, which can be summarized as a spectral hypergraph k-way partitioning problem [20]. And the normalized one cut spectral hypergraph partitioning task is defined as follows:
@@
Assuming a key-way hypergraph partition by a set of vertexes subsets $\{V_{1},\cdots,V_{p}\}$ from $\mathbf{F}=[f_{1}\cdots f_{p}]$. And then the spectral hypergraph k-way partitioning problem as a combinatorial optimization problem can be relaxed with respect to minimizing $\beta(V_{1},\cdots,V_{p})$ as follows:
$\mathop{\arg\min}\limits_{\beta}\{\beta(V_{1},\cdots,V_{p})\}=\{\sum_{i=1}^{p}f_{i}^{\top}\Delta f_{i}\}=\{trace(\mathbf{F}^{\top}\Delta\mathbf{F})\}$ (7)
Firstly, we utilize the K-nearest neighbor (KNN) method [21, 22] with a hyperparameter $k$ to generate a set of hypergraphs which presents group-wise interactions of crowded movements. Suppose the trajectory feature of $i$th agent $\mathbf{q}({x}_{i})\in{\mathbb{R}}^{d}$ and the similarity matrix of crowded movements $\mathcal{S}\in{\mathbb{R}}^{N\times N}$. We utilize a fully connected layer (FC) to capture the low-level trajectory embeddings of each agent as $\{\mathbf{q}_{(\rm{X}_{\emph{1}})},\cdots,\mathbf{q}_{(\rm{X}_{\emph{N}})}\}$ from input $\{\mathbf{X}_{\emph{1}},\cdots,\mathbf{X}_{\emph{N}}\}$. And then the Mahalanobis Distance [23] as a modified Euclidean Distance is used to calculate agent pairs’ feature distances and their similarity matrix. Considering the property of distributions, we construct the social interaction hypergraph based on feature dimensional Mahalanobis Distance, in which avoids the potential errors of Euclidean Distance. Apart from that, Mahalanobis Distance is also deployed to reduce the interferences from feature correlations and measurement dimensions. The Mahalanobis Distance of $(i,j)$-th vertex pair $Dis(i,j)$ in the feature dimension is defined as follows:
@@
The $(i,j)$ element of similarity matrix $\mathcal{S}$ is defined as follows:
@@
where $\varrho$ is the median value of all vertex pairs’ feature distances.
Once obtaining the similarity matrix $\mathcal{S}$, KNN constructs multiscale hypergraphs based on feature Mahalanobis Distance as shown in Fig. 3. Each vertex of hypergraph represents singl agent and each hyperedge could link multiple agents to capture group-wise interactions. Especially, herein each hyperedge is created by linking one vertex and its $k$ nearest neighbors with respect to similarity matrix $\mathcal{S}$.
@@
Figure 4: Hybrid Spatial-Temporal Transformer Framework: Pedestrians’ motion intents and dependencies are abstracted as spatial and temporal attention maps by multi-head attention mechanism of spatial-temporal transformer. Additionally, a multi-head cross attention mechanism is employed to align heterogeneous spatial-temporal features.
@@
III-E Multi-Modal Transformer and CVAE Decoder
A multi-modal transformer is developed to fuse heterogeneous group-wise and pair-wise agent interaction features based on [15], using a cross-model attention layer. Wherein the multihead cross-modal attention mechanism $\rm CMAtten(\cdot)$ aligns each unit modality with the concatenation of all modalities $\mathbf{Y}_{\rm F}=Concate(\mathbf{\hat{Y}}_{\rm H},\mathbf{\hat{Y}}_{\rm S},\mathbf{\hat{Y}}_{\rm T})$ as follows:
$\rm{CMAtten}(\hat{Y}_{U})$ $=\rm{Multi}(\mathbf{Q}^{\rm{head_{j}}}_{\rm{U}},\mathbf{K}^{\rm{head_{j}}}_{\rm{F}},\mathbf{V}^{\rm{head_{j}}}_{\rm{F}})$ (15)
where $\rm U\in\{\rm{S},\rm{T},\rm{H}\}$, and $\mathbf{Y}^{\rm{head_{j}}}_{\rm{U}}$ present the $\rm j$-th head crossmodal attention of arbitrary unit modal with $\rm j\in\{1,\cdots,h\}$, and $\mathbf{Q}^{\rm{head_{j}}}_{\rm{U}}\in\mathbb{R}^{\rm{L}_{U}\times\rm{d}_{k}},\mathbf{K}^{\rm{head_{j}}}_{\rm{F}}\in\mathbb{R}^{\rm{L}_{F}\times\rm{d}_{k}},\mathbf{V}^{\rm{head_{j}}}_{\rm{F}}\in\mathbb{R}^{\rm{L}_{F}\times\rm{d}_{k}}$ are unit modality query and fused modality key or value.
In the multi-modal transformer block as Fig. 4, group-wise feature $\mathbf{\hat{Y}_{{H}}}$ and pair-wise feature $\{\mathbf{\hat{Y}_{{S}}},\mathbf{\hat{Y}_{{T}}}\}$ are fed into a positional embedding layer separately, in which all unit modality are concatenated as fused modality $\mathbf{Y}_{\rm F}$. And then the multi-head cross-modal attention block is utilized to gather above group-wise and pair-wise features. Lastly, three fused high-level embeddings ($\mathbf{\hat{Y}_{\rm{H}}}\&\mathbf{{Y}_{\rm{F}}},\mathbf{\hat{Y}_{\rm{S}}}\&\mathbf{{Y}_{\rm{F}}},\mathbf{\hat{Y}_{\rm{T}}}\&\mathbf{{Y}_{\rm{F}}}$) are processed by a multi-head self-attention layer as final state representation of crowd environmental dynamics.

### v2

@@ Abstract
Predicting crowded intents and trajectories is crucial in various real-world applications, including social robots and autonomous vehicles. Understanding environmental dynamics is challenging, not only due to the complexities of modeling pair-wise spatial and temporal interactions but also the diverse influence of group-wise interactions. To decode the comprehensive pair-wise and group-wise interactions in crowded scenarios, we introduce Hyper-STTN, a Hypergraph-based Spatial-Temporal Transformer Network for crowd trajectory prediction. In Hyper-STTN, crowded group-wise correlations are constructed using a set of multi-scale hypergraphs with varying group sizes, captured through random-walk probability hypergraph spectral convolution. Additionally, a spatial-temporal transformer is adapted to capture pedestrians’ pair-wise latent interactions in spatial-temporal dimensions. These heterogeneous group-wise and pair-wise are then fused and aligned though a multi-modal transformer network. Hyper-STTN outperforms other state-of-the-art baselines and ablation models on several public pedestrian motion datasets.
@@ I Introduction
Human trajectory prediction is a pivotal area in computer vision and robotics research, focusing on predicting agents’ future movements by analyzing their past behaviors. This is vital for many real-world applications such as smart city systems and social robot navigation [1, 2]. Despite its importance, forecasting human trajectories in social settings is complex due to the unpredictable nature and varied patterns of crowd movements. The challenge is intensified by the need to understand the intricate interactions and relations among different social groups, which involve complex cooperation and competitive dynamics, as shown in Fig. 1.
@@
Figure 1: HHI (human-human interaction) illustration: Group-wise HHI captures latent correlations among high-level perspectives on group behaviors, while pair-wise spatial-temporal HHI represents individual influences.
@@
To address these challenges, we introduce Hyper-STTN, a hypergraph-based spatial-temporal transformer network designed to handle both pair-wise and group-wise underlying social interactions, as illustrated in Fig. 1, across spatial-temporal dimensions. The core concept behind Hyper-STTN is its utilization of an effective HHI and social interaction inference framework using multiscale hypergraphs. The main contributions of this paper can be summarized as follows: (1). We propose a novel human trajectory prediction framework called Hyper-STTN, which constructs crowded dynamics using a set of multiscale hypergraphs to leverage group-wise and pair-wise social interactions in both spatial and temporal dimensions; (2). Hyper-STTN aligns heterogeneous features and interactions using a multi-modal transformer network to avoid confusion in interpretations; (3). Hyper-STTN demonstrates outstanding performance compared with other existing SOTA approaches on publicly available pedestrian trajectory datasets.
@@ II-A Related Works
Earlier efforts formulated the task of predicting human trajectories based on handcrafted models, such as the social force model [10] and Gaussian process [11], which utilized fixed physical or mathematical rules to adapt to environmental dynamics. The inflexibility of traditional approaches could lead to overlooking underlying social interaction and forecasting errors in more complex conditions. Inspired by the advancements in machine learning, many deep learning-based algorithms have been deployed to better understand social interactions among pedestrians [4, 6]. However, previous learning-based methods either could not fully utilize the captured features directly or could not effectively leverage long-term dependencies. Especially, due to the successes of attention mechanisms and transformers in sequence learning and pair-wise feature representation [12, 7, 13], there have been explicit descriptions of social interactions via both spatial and temporal dimensions using a spatio-temporal graph model of crowd movements. While the description of pair-wise interactions has been well addressed with the development of transformers and attention modules, all the aforementioned methods have neglected to collect group-wise latent influences for individuals.
While [5] has adapted a hypergraph-based network to learn group-wise and pair-wise features with the multiscale property of hypergraph generation, GroupNet still struggles with isolating spatial features along the temporal dimension. Therefore, the challenge of aligning both group-wise and pair-wise social interactions with long-term dependencies remains a key problem in human trajectory prediction. In this work, we not only incorporate spatial and temporal pair-wise interactions into hypergraph-based group-wise interaction representation, but also introduce a cross-modal attention mechanism to leverage the heterogeneous multi-modality dependencies, based on a multi-modal transformer [14].
@@ III-A Preliminary
We adapt a set of multiscale hypergraphs $\mathcal{G}=(\mathcal{V,E,T}\mathbf{,W})$ to construct crowd movement dynamics for trajectory prediction inference. Where $\mathcal{V}=\{{v_{\rm 1}},{v_{\rm 2}},\cdots,{v_{\rm N}}\}$ is the set of vertexes that presents total ${\rm N}$ agents in the scenario, and =$\mathcal{E}=\{{e_{\rm 1}},{e_{\rm 2}},\cdots,{e_{\rm M}}\}$ denotes that a set of ${\rm M}$ hyperedges are constructed in the hypergraphs, particularly each hyperedge can link more than two vertexes to form a group, and $\mathcal{T}=\{{\tau_{\rm 1}},{\tau_{\rm 2}},\cdots,{\tau_{\rm H}}\}$ presents $\rm H$ scales of hypergrpahs, and $\mathbf{W}=diag({{w}_{\rm e_{1}}},{{w}_{\rm e_{2}}},\cdots,{{w}_{\rm e_{M}}})\in\mathbb{R}^{M\times M}$ defines a weighted matrix of hypergraphs. Lastly, the incident matrix $\mathbf{H}\in\mathbb{R}^{N\times M}$ of hypergraph $\mathcal{G}$ is defined as follows: $\mathbf{H}(v,e)=1,~{}\text{if}~{}v\in e;~{}\mathbf{H}(v,e)=0,~{}\text{if}~{}v\notin e.$ The diagonal vertex matrix $\mathbf{M}_{v}\in\mathbb{R}^{N\times N}$ and diagonal edge matrix $\mathbf{M}_{e}\in\mathbb{R}^{M\times M}$ are composed by the degree of vertex $d(v)=\sum_{e\in\mathcal{E}}w_{(e)}\mathbf{H}(v,e)$ and the degree of edge $d(e)=\sum_{v\in\mathcal{V}}\mathbf{H}(v,e)$.
Additionally, the adjacency matrix of hypergraph $\mathcal{G}$ can be represented by $\mathbf{A}\in\mathbb{R}^{N\times N}$, whereas $\mathbf{A}=\mathbf{H}\mathbf{W}\mathbf{H}^{\top}-\mathbf{M}_{\rm v}$. Hyper-STTN not only models group-wise interaction using pedestrians hypergraph construction, but also leverages individual spatial and temporal interaction to align both group-wise and pair-wise features. Let $\mathbf{X}\in\mathbb{R}^{N\times T_{i}\times d}$ and $\mathbf{\hat{X}}\in\mathbb{R}^{N\times T_{o}\times d}$ denote the past and future trajectory data of crowd with time sequence $T_{i}=8$ or $T_{o}=12$ and dimension $d=2$. And the position of $i$th agent in time $t$ can be defined as $\mathbf{x}_{i}^{t}=(x,y)$. Lastly, Hyper-STTN describes the pedestrian trajectory distribution $\mathcal{P}$ as follows:
@@ III-B Hyper-STTN Architecture
Hyper-STTN captures both group-wise and pair-wise interactions to reason out HHI for human trajectory prediction tasks, as illustrated in Fig. 2. Firstly, the spatial-temporal transformer network is developed to infer spatial and temporal dimensional pair-wise features. Meanwhile, the hypergraph convolution neural network is employed in parallel to present crowd group-level dependencies, based on crowded multiscale hypergraph construction. Subsequently, above heterogeneous group-wise and pair-wise crowded interaction matrices are fused through a multi-modal transformer to align the spatial-temporal embeddings into the hypergraph. Finally, the crowd embedding mixture is decoded from a CVAE [18] decoder to forecast future trajectories.
@@ III-C Spatial-Temporal Transformer
Inspired by [19, 7, 1], we have developed a spatial-temporal transformer network to abstract the correlation of agents’ movement along spatial-temporal dimensions. Particularly, spatial or temporal dependencies can be represented as multiple attention maps, as shown in Fig. 6. In the spatial transformer, the input data $\mathbf{X}$ is first fed into a positional encoding (PE) layer for sequence information encoding based on [12]. Subsequently, the input data proceeds in the proper order through a layer normalization block (LN), a multihead spatial attention layer, and a feed-forward network (FFN), as shown in Fig. 4. Additionally, residual connections are deployed to stabilize the network. The core of the spatial-temporal transformer is the attention mechanism, where the calculation of the multihead spatial or temporal attention layer is defined as follows:
@@
where $f_{\rm{fc}}$ is a fully connected layer, $\rm{head}_{(\cdot)}$ denotes $\rm i$-th head in the multihead attention with $\rm{i}\in[1,\cdots,\rm{h}]$, $\rm\mathbf{Q}^{i},\rm\mathbf{K}^{i},\rm\mathbf{V}^{i}$ are $\rm i$-th query matrix, key matrix, and value matrix with the dimension $\rm d_{h}$.
Figure 3: Group-wise HHI Representation: i) We construct group-wise HHI with a set of multiscale hypergraphs, where each agent is queried in the feature space with varying ’k’ in KNN to link multiscale hyperedges. ii) After constructing HHI hypergraphs, group-wise dependencies are captured by point-to-edge and edge-to-point phases with hypergraph spectral convolution operations.
@@
The framework of temporal transformer and spatial transformer are illustrated as Fig. 4. The spatial transformer and temporal transformer capture pair-wise spatial interaction $\mathbf{\hat{Y}}_{\rm{S}}\in\mathbb{R}^{N\times T_{i}\times d}$ along each timestep and individual’s temporal interaction $\mathbf{\hat{Y}}_{\rm{T}}\in\mathbb{R}^{N\times T_{i}\times d}$ along each agent.
@@ III-D Hypergraph Neural Network
The crowd groups of Hyper-STTN are constructed with respect to neighborhoods’ feature dimensional Mahalonobis distance [21], considering both interactive correlation distribution and physical spatial distance. In particular, the pedestrians’ motion similarity is captured by the covariance matrix of Mahalonobis distance to cover the situations that a long-distance person presents a high potentiality to join the group. We construct crowd hypergraphs by vertex classifications of social group which are formulated as a spectral hypergraph k-way partitioning problem [22, 23]. The normalized one cut spectral hypergraph partitioning task is defined as follows:
@@
Assuming the $k$-way hypergraph partition by a set of vertexes subsets $\{\mathcal{V}_{1},\cdots,\mathcal{V}_{p}\}$ from $\mathbf{F}=[f_{1}\cdots f_{p}]$. Subsequently, the spectral hypergraph $k$-way partitioning problem as a combinatorial optimization problem can be relaxed with respect to minimizing $\beta(\mathcal{V}_{1},\cdots,\mathcal{V}_{p})$ as follows:
$\mathop{\arg\min}\limits_{\beta}\{\beta(\mathcal{V}_{1},\cdots,\mathcal{V}_{p})\}=\{\sum_{i=1}^{p}f_{i}^{\top}\Delta f_{i}\}=\{trace(\mathbf{F}^{\top}\Delta\mathbf{F})\}$ (7)
To address above k-way hypergraph partition task, the K-nearest neighbor (KNN) [24] method is leveraged for multiscale hypergraphs generation. Hyper-STTN iterates each vertex who presents a person with KNN algorithm to search several its interactive neighborhoods based on Mahalonobis feature distance. In details, the low-level trajectory embeddings of each agent $\{\mathbf{q}_{(\rm{X}_{\emph{1}})},\cdots,\mathbf{q}_{(\rm{X}_{\emph{N}})};\mathbf{q}({x}_{i})\in{\mathbb{R}}^{d}\}$ is encoded by a fully connected layer (FC) from input $\{\mathbf{X}_{\emph{1}},\cdots,\mathbf{X}_{\emph{N}}\}$. Subsequently, the similarity matrix $\mathcal{S}$ is filled by feature dimensional agent pairs’ Mahalanobis distance [21], coupling the interaction attributes and motion distributions to dilute Euclidean distance analogous interferences and errors from feature correlations and distributions. The Mahalanobis distance of $(i,j)$-th vertex pair $Dis(i,j)$ in the feature dimension is defined as follows:
@@
The $(i,j)$ element of similarity matrix $\mathcal{S}\in{\mathbb{R}}^{N\times N}$ is defined as follows:
@@
where $\varrho$ is the mean of all vertex pairs’ feature distances.
Inspired by [25, 23], KNN searches each vertex and its $K$ nearest neighbors to construct multiscale hypergraphs with respect to the similarity matrix $\mathcal{S}$, as shown in Fig. 3. Each vertex of hypergraph represents single agent, and hyperedges could link multiple agents to form group-wise interactions.
@@
Figure 4: Hybrid Spatial-Temporal Transformer Framework: Pedestrians’ motion intents and dependencies are abstracted as spatial and temporal attention maps by multi-head attention mechanism of spatial-temporal transformer. Additionally, a multi-head cross attention mechanism is employed to align heterogeneous group-wise and pair-wise features.
@@
III-E Multi-Modal Fusion Network and CVAE Decoder
As shown in Fig. 4, the multi-modal transformer [14] is developed to fuse heterogeneous spatial-temporal features, using a cross-attention layer and self-attention layer. Wherein the multihead cross-attention mechanism $\rm CMAtten(\cdot)$ captures cross-modality features as follows:
$\rm{CMAtten}(\hat{Y}_{U})$ $=\rm{Multi}(\mathbf{Q}^{\rm{head_{j}}}_{\rm{U}},\mathbf{K}^{\rm{head_{j}}}_{\rm{U^{\prime}}},\mathbf{V}^{\rm{head_{j}}}_{\rm{U^{\prime}}})$ (15)
where $\rm U\in\{\rm{S},\rm{T},\rm{H}\}$, and $\mathbf{Y}^{\rm{head_{j}}}_{\rm{U}}$ present the $\rm j$-th head crossmodal attention of arbitrary unit modal with $\rm j\in\{1,\cdots,h\}$.
Subsequently, the group-wise HHI features $\mathbf{\hat{Y}}_{\rm{HS}},\mathbf{\hat{Y}}_{\rm{HT}}$ and pair-wise HHI dependencies $\mathbf{\hat{Y}}_{\rm{ST}}$ are aligned by a multi-head self-attention network as final crowd dynamics representation $\mathbf{\hat{Y}}_{\rm{M}}$.